# MindCraft: How Concept Trees Take Shape In Deep Models

### v1

@@ 1 Introduction
Deep learning has achieved remarkable success across diverse domains, including computer vision (Krizhevsky et al., 2012), text generation (Devlin et al., 2019), and speech recognition (Hinton et al., 2012; Graves et al., 2013). However, the internal mechanisms of neural networks remain opaque. Despite rapid progress in visualization and interpretability techniques, we still lack a clear understanding of how abstract concepts, such as “disease,” “cause,” or “truth”, form and stabilize inside their layers (Lipton, 2016; Doshi-Velez & Kim, 2017b; Ribeiro et al., 2016). This opacity has fueled the widespread description of neural networks as “black boxes” Lipton (2016); Doshi-Velez & Kim (2017b), raising concerns about reliability in sensitive domains such as healthcare (Caruana et al., 2015), finance (Rudin, 2019), and law Doshi-Velez & Kim (2017a).
Recent advances in representation analysis have opened promising avenues for transparency. For example, networks trained to play chess acquired a range of human chess concepts (McGrath et al., 2022). Similarly, generative and self-supervised models exhibit emergent representations such as semantic segmentation in vision tasks (Caron et al., 2021; Oquab et al., 2023). Zou et al. (2023) formalized this line of work as Representation Engineering (RepE), showing how we can extract concept directions from a model’s internal states and even control its behavior. RepE provides a top-down view of interpretability, shifting focus from individual neurons or circuits to global representational structure. In parallel, the Linear Representation Hypothesis (LRH) (Park et al., 2023) suggests that task-relevant concepts gradually become linearly separable while the input propagates through deeper layers of the model, enabling simple linear probes and edits to access them.
@@ 2.1 Representation Learning in Deep Networks
Representation Learning studies how neural networks develop internal representations that encode task-relevant information. Early research on word embeddings demonstrated that neural networks can capture rich semantic and syntactic relationships in a distributed manner (Mikolov et al., 2013). Subsequent studies showed that such representations often reflect abstract latent factors: for example, Radford et al. (2018) observed the emergence of sentiment-tracking units, and Schramowski et al. (2019) reported that large language models (LLMs) encode implicit moral dimensions.
Similar phenomena have been observed beyond language. In reinforcement learning, McGrath et al. (2022) found that models trained to play chess develop internal abstractions of board state and strategy. In computer vision, self-supervised and generative objectives have been shown to induce semantic feature maps useful for downstream tasks such as segmentation (Caron et al., 2021; Oquab et al., 2023). Building on these observations, Zou et al. (2023) introduced Representation Engineering (RepE), which uses contrastive edits to extract and manipulate concept directions from internal activations. Their work exemplifies a top-down approach to interpretability: instead of focusing on single neurons or weights, it analyzes the global structure of representations. Similarly, Park et al. (2023) proposed the Linear Representation Hypothesis (LRH), arguing that task-relevant concepts become linearly separable with depth, enabling simple linear probes and edits.
@@ 2.2 Neural Network Interpretability
Many interpretability methods emphasize parameter or gradient-based structures, seeking to explain model behavior in terms of weights, activations, or neuron circuits. Saliency-based methods including Simonyan et al. (2013); Springenberg et al. (2014); Zhou et al. (2016) highlight influential input regions via gradients or activations, while feature visualization (Zeiler & Fergus, 2014) synthesizes inputs that strongly activate particular neurons. More recently, mechanistic interpretability aims to reverse engineer networks into circuits of interpretable components (Olah et al., 2020; Olsson et al., 2022; Lieberum et al., 2023). Although powerful, these approaches can be brittle or require extensive manual effort, and they typically do not explain how abstract representations form from distributed activations.
Recent works instead analyze the geometry and dynamics of representation spaces. Techniques such as CKA (Kornblith et al., 2019), SVCCA (Raghu et al., 2017), and Procrustes alignment (Gao et al., 2021) measure how representations evolve across layers or between models. Other studies have investigated how residual connections promote stable feature reuse and facilitate optimization (He et al., 2016; Dehghani et al., 2023), and how singular value spectra reflect information propagation and compression in deep networks (Saxe et al., 2014; Morcos et al., 2018; Canatar et al., 2021). These lines of work suggest that deep networks may progressively concentrate representational energy along a few dominant directions, a phenomenon that we directly exploit.
@@ 3.1 Counterfactuals
We adopt the notion of counterfactual from causal inference (Pearl, 2009), defined as the resulting variable after an intervention is applied to a particular variable (e.g., a token), while the surrounding context is held constant. The resulting variable, observed after the intervention, is the counterfactual to the original variable. Specifically, if we intervene on a specific component $x$ of $X$, and calculate the counterfactual on the general target variable $Y$, we write:
@@ 4.1 Motivation: Propagation of Counterfactual Signals
The only difference is the word change (powerful $\rightarrow$ powerless). By analyzing how this minimal intervention propagates through the network (see Figure 1), we can observe where the model begins to treat the two inputs differently. Following prior work showing that the last token often best captures a model’s generative state (Meng et al., 2022; Zou et al., 2023), we focus on the value representations of the final token in the sequence, denoted as $V_{L(\Delta x)}^{(-1)}$ and $V_{L}^{(-1)}$, where the superscript 1-1 indicates the last token in the sequence. Comparing these across layers reveals how the counterfactual difference evolves.
@@
Figure 2 shows the cosine similarity between the factual and counterfactual representations at each layer. In the early layers (roughly before layer 10), similarity remains high, meaning the network has not yet distinguished the two contexts. In mid-layers, however, similarity drops sharply (red bars), indicating a sudden amplification of the conceptual difference. Beyond these branching points, the representations stabilize again, converging toward a consistent trajectory in deeper layers.
Figure 2: $\cos(V_{L(\Delta x)}^{(-1)},V_{L}^{(-1)})$ change.
@@ 4.2 Paving the Concept Path
Our analysis focuses on the last token in the sequence, since it integrates information from the full context and directly conditions the model’s next prediction  (Meng et al., 2022; Zou et al., 2023). By examining how this representation changes across layers, we can trace the emergence of conceptual differences. Therefore, we focus on the last-token representation, denoted $Z^{(-1)}\in\mathbb{R}^{d}$, as the anchor point for our spectral analysis. While the Value vector of the last token already carries rich contextual information, directly analyzing it can be noisy and unstable. To obtain a more robust basis, we examine the Value transformation matrix $W_{V}$ using singular value decomposition (SVD):
@@
where $U\Sigma R^{\top}$ is its SVD with $U=[u_{1},\dots,u_{m}]$, $\Sigma=\mathrm{diag}(\sigma_{1},\dots,\sigma_{p})$, $R=[r_{1},\dots,r_{n}]$, and $p=\min(m,n)$. These left singular vectors, $u_{i}$, represent the principal directions along which the transformation $W_{V}$ has the most significant amplifying or attenuating effect, as determined by the corresponding singular values in $\Sigma$. This basis provides a more stable and meaningful coordinate system to analyze than the content of the raw Value vector, which is further testified in Appendix A. With this setup, we define the Concept Path as the decomposition of the last token’s Value vector across all principal directions of $W_{V}$. This gives us a spectral signature of the token’s information content.
@@ 5.1 Concept Tree Implementations
Figure 3: Concept Trees constructed from six scenarios: (a–b) medical diagnosis, (c) daily life, (d) personality evaluation, (e) physics reasoning, and (f) political decision-making. Counterfactual tokens are marked with underlines, and n denotes the number of remaining unbranched concepts.
The experiment results from a variety of scenarios shown in Figure 3. For concept identification, we incorporate the automated pipeline as demonstrated in Appendix B. We use $k=10$ and $\tau=0.9$ by default, and more details about experiment settings are in Appendix C. The results demonstrate the broad applicability of the concept tree, and we find that the model’s internal processing of concepts is indeed not a monolithic, linear process but rather a structured, hierarchical tree-shaped organization. This tree serves as a powerful visualization, mapping the pathway of concept transformations—from the most influential conceptual directions to progressively less contributive ones—that the model follows to derive a conclusion, such as reasoning from one fact to another.
@@ 5.2 Concept Highlight Experiment
To further validate that the Concept Tree framework captures semantic interpretability rather than relying solely on static measures such as specific embeddings, we perform an experiment in Figure 4. In case (h), the model processes a standard sentence without explicit emphasis, where temporal tokens such as “2024” and “2025” behave similarly to other factual tokens. In contrast, case (g) introduces an explicit emphasis on the temporal information by appending the instruction “in this sentence, 2024 or 2025 is extremely important”. Consequently, this analysis shows that case (h) downplays the temporal tokens, while case (g) induces an earlier branching that clearly separates “2024” and “2025”. Therefore, this modification significantly increases the conceptual salience of the “2024/2025” pair, elevating their importance in the reasoning process. In conclusion, this experiment highlights that our method captures conceptual importance shaped by context, rather than mechanical token embedding distances.
Figure 4: Concept Highlight Experiment.
Figure 5: Layer-wise analysis of attention weights, value vectors, and representations, where Counterfactual Pair 1 is “Pretend you’re an honest (untruthful) person making statements about the world.”, Counterfactual Pair 2 is “ Describe a fair (biased) scenario that you have seen.”, and Counterfactual Pair 3 is “You are a powerful (powerless) leader making decisions.”, we provide average and standard deviation of five random perturbations of the original context.
@@ 5.4 Disentanglement of Input Embeddings and Concepts
Another central question is whether the model’s separation of high-level concepts is statistically dependent on the input embedding distances. To investigate this, we analyze the relationship between the initial distance of input token embeddings and the layer at which their corresponding concepts first diverge in the Concept Tree framework. We quantify the embedding distance using the L2 distance, shown in Figure 6. The scatter plot in Figure 6 (a) does not reveal a clear linear relationship between the L2 distance and the Branching Layer. In other words, tokens that are farther apart in embedding space do not consistently separate earlier in the network. This observation is further supported by the correlation analysis in Figure 6 (b), where the overall Pearson’s r=0.218r=-0.218 and Spearman’s $\rho=-0.102$ remain statistically insignificant across most cases. While certain contexts, such as case (c), show a stronger negative correlation, the variability across different cases suggests that token-level embedding distances alone cannot fully explain when concepts diverge. Taken together, these results indicate the disentanglement between low-level token embeddings and high-level conceptual organization. This, in turn, highlights the flexibility and broad potential of the Concept Tree framework for extracting and analyzing abstract concepts beyond what is encoded in static embeddings.
@@
Figure 6: Comparison of L2 distance against separation layer and their correlation analysis.
@@ 6 Conclusion
We employ Qwen2.5-7B-Instruct (Team, 2024), a 7-billion-parameter instruction-tuned large language model developed by Alibaba’s Qwen team as part of the Qwen2.5 series. It adopts a decoder-only Transformer architecture with improvements such as SwiGLU activations and group query attention, and supports very long context lengths of up to 128K tokens. All experiments can be run on a single NVIDIA A100 GPU.
@@
The core of our methodology is the construction of Concept Trees from counterfactual pairs. Key hyperparameters, such as the separation threshold ($\tau=0.9$) and the number of top-$k$ components ($k=10$) for the Conceptual Separation Score, are specified in our code and the corresponding sections. The six scenarios used for our main experiments (Figure 7) are detailed in the appendix, along with the specific counterfactual pairs used for each. The automated pipeline for generating new counterfactuals is described in Appendix B.
Appendix A The Comparison between Raw Value Matrix and SVD

### v2

@@ 1 Introduction
Deep learning has achieved remarkable success across diverse domains, including computer vision (krizhevsky2012imagenet), text generation (devlin2019bert), and speech recognition (hinton2012deep; graves2013speech). However, the internal mechanisms of neural networks remain opaque. Despite rapid progress in visualization and interpretability techniques, we still lack a clear understanding of how abstract concepts, such as “disease,” “cause,” or “truth”, form and stabilize inside their layers (lipton2016mythos; doshi2017towards; ribeiro2016should). This opacity has fueled the widespread description of neural networks as “black boxes” lipton2016mythos; doshi2017towards, raising concerns about reliability in sensitive domains such as healthcare (caruana2015intelligible), finance (rudin2019stop), and law doshi2017towards.
Recent advances in representation analysis have opened promising avenues for transparency. For example, networks trained to play chess acquired a range of human chess concepts (mcgrath2022acquisition). Similarly, generative and self-supervised models exhibit emergent representations such as semantic segmentation in vision tasks (caron2021emerging; oquab2023dinov2). zou2023representation formalized this line of work as Representation Engineering (RepE), showing how we can extract concept directions from a model’s internal states and even control its behavior. RepE provides a top-down view of interpretability, shifting focus from individual neurons or circuits to global representational structure. In parallel, the Linear Representation Hypothesis (LRH) (park2023linear) suggests that task-relevant concepts gradually become linearly separable while the input propagates through deeper layers of the model, enabling simple linear probes and edits to access them.
@@ 2.1 Representation Learning in Deep Networks
Representation Learning studies how neural networks develop internal representations that encode task-relevant information. Early research on word embeddings demonstrated that neural networks can capture rich semantic and syntactic relationships in a distributed manner (mikolov2013distributed). Subsequent studies showed that such representations often reflect abstract latent factors: for example, radford2018learning observed the emergence of sentiment-tracking units, and schramowski2019bert reported that large language models (LLMs) encode implicit moral dimensions.
Similar phenomena have been observed beyond language. In reinforcement learning, mcgrath2022acquisition found that models trained to play chess develop internal abstractions of board state and strategy. In computer vision, self-supervised and generative objectives have been shown to induce semantic feature maps useful for downstream tasks such as segmentation (caron2021emerging; oquab2023dinov2). Building on these observations, zou2023representation introduced Representation Engineering (RepE), which uses contrastive edits to extract and manipulate concept directions from internal activations. Their work exemplifies a top-down approach to interpretability: instead of focusing on single neurons or weights, it analyzes the global structure of representations. Similarly, park2023linear proposed the Linear Representation Hypothesis (LRH), arguing that task-relevant concepts become linearly separable with depth, enabling simple linear probes and edits.
@@ 2.2 Neural Network Interpretability
Many interpretability methods emphasize parameter or gradient-based structures, seeking to explain model behavior in terms of weights, activations, or neuron circuits. Saliency-based methods including simonyan2013deep; springenberg2014striving; zhou2016learning highlight influential input regions via gradients or activations, while feature visualization (zeiler2014visualizing) synthesizes inputs that strongly activate particular neurons. More recently, mechanistic interpretability aims to reverse engineer networks into circuits of interpretable components (olah2020zoom; olsson2022context; lieberum2023does). Although powerful, these approaches can be brittle or require extensive manual effort, and they typically do not explain how abstract representations form from distributed activations.
Recent works instead analyze the geometry and dynamics of representation spaces. Techniques such as CKA (kornblith2019similarity), SVCCA (raghu2017svcca), and Procrustes alignment (gao2021simcse) measure how representations evolve across layers or between models. Other studies have investigated how residual connections promote stable feature reuse and facilitate optimization (he2016deep; dehghani2023scaling), and how singular value spectra reflect information propagation and compression in deep networks (saxe2013exact; morcos2018importance; canatar2021spectral). These lines of work suggest that deep networks may progressively concentrate representational energy along a few dominant directions, a phenomenon that we directly exploit.
@@ 3.1 Counterfactuals
We adopt the notion of counterfactual from causal inference (pearl2009causality), defined as the resulting variable after an intervention is applied to a particular variable (e.g., a token), while the surrounding context is held constant. The resulting variable, observed after the intervention, is the counterfactual to the original variable. Specifically, if we intervene on a specific component $x$ of $X$, and calculate the counterfactual on the general target variable $Y$, we write:
@@ 4.1 Motivation: Propagation of Counterfactual Signals
The only difference is the word change (powerful $\rightarrow$ powerless). By analyzing how this minimal intervention propagates through the network (see Figure 1), we can observe where the model begins to treat the two inputs differently. Following prior work showing that the last token often best captures a model’s generative state (meng2022locating; zou2023representation), we focus on the value representations of the final token in the sequence, denoted as $V_{L(\Delta x)}^{(-1)}$ and $V_{L}^{(-1)}$, where the superscript 1-1 indicates the last token in the sequence. Comparing these across layers reveals how the counterfactual difference evolves.
@@
Figure 2 shows the cosine similarity between the factual and counterfactual representations at each layer. In the early layers (roughly before layer 10), similarity remains high, meaning the network has not yet distinguished the two contexts. In mid-layers, however, similarity drops sharply (red bars), indicating a sudden amplification of the conceptual difference. Beyond these branching points, the
Figure 2: $\cos(V_{L(\Delta x)}^{(-1)},V_{L}^{(-1)})$ change, where a sudden amplification of the conceptual difference is observed.
representations stabilize again, converging toward a consistent trajectory in deeper layers.
@@ 4.2 Paving the Concept Path
Our analysis focuses on the last token in the sequence, since it integrates information from the full context and directly conditions the model’s next prediction  (meng2022locating; zou2023representation). By examining how this representation changes across layers, we can trace the emergence of conceptual differences. Therefore, we focus on the last-token representation, denoted $Z^{(-1)}\in\mathbb{R}^{d}$, as the anchor point for our spectral analysis. While the Value vector of the last token already carries rich contextual information, directly analyzing it can be noisy and unstable. To obtain a more robust basis, we examine the Value transformation matrix $W_{V}$ using singular value decomposition (SVD):
@@
where $U\Sigma R^{\top}$ is its SVD with $U=[u_{1},\dots,u_{m}]$, $\Sigma=\mathrm{diag}(\sigma_{1},\dots,\sigma_{p})$, $R=[r_{1},\dots,r_{n}]$, and $p=\min(m,n)$. These left singular vectors, $u_{i}$, represent the principal directions along which the transformation $W_{V}$ has the most significant amplifying or attenuating effect, as determined by the corresponding singular values in $\Sigma$. This basis provides a more stable and meaningful coordinate system to analyze than the content of the raw Value vector, which is further testified in Appendix B.1. With this setup, we define the Concept Path as the decomposition of the last token’s Value vector across all principal directions of $W_{V}$. This gives us a spectral signature of the token’s information content.
@@ 5.1 Concept Tree Implementations
Figure 3: Concept Trees constructed from six scenarios: (a–b) medical diagnosis, (c) daily life, (d) personality evaluation, (e) physics reasoning, and (f) political decision-making. Counterfactual tokens are marked with underlines, and n denotes the number of remaining unbranched concepts. The results demonstrate the broad applicability of the Concept Tree, revealing a structured and hierarchical organization of conceptual reasoning within the model.
The experiment results from a variety of scenarios shown in Figure 3. For concept identification, we incorporate the automated pipeline as demonstrated in Appendix C. We use $k=10$ and $\tau=0.9$ by default, and more details about experiment settings are in Appendix D. The results demonstrate the broad applicability of the concept tree, and we find that the model’s internal processing of concepts is indeed not a monolithic, linear process but rather a structured, hierarchical tree-shaped organization. This tree serves as a powerful visualization, mapping the pathway of concept transformations—from the most influential conceptual directions to progressively less contributive ones—that the model follows to derive a conclusion, such as reasoning from one fact to another.
@@ 5.2 Concept Highlight Experiment
To further validate that the Concept Tree framework captures semantic interpretability rather than relying solely on static measures such as input or latent embeddings, we perform an experiment in Figure 4. In case (h), the model processes a standard sentence without explicit emphasis, where temporal tokens such as “2024” and “2025” behave similarly to other factual tokens. In contrast, case (g) introduces an explicit emphasis on the temporal information by appending the instruction “in this sentence, 2024 or 2025 is extremely important”. Consequently, this analysis shows that case (h) downplays the temporal tokens, while case (g) induces an earlier branching that clearly separates “2024” and “2025”. Therefore, this modification significantly increases the conceptual salience of the “2024/2025” pair, elevating their importance in the reasoning process. In conclusion, this experiment highlights that our method captures conceptual importance shaped by context, rather than mechanical token embedding distances.
Figure 4: Concept Highlight Experiment. This experiment highlights Concept Tree captures semantic interpretability rather than relying solely on static measures such as input or latent embeddings.
Figure 5: Layer-wise analysis of attention weights, value vectors, and representations, where Counterfactual Pair 1 is “Pretend you’re an honest (untruthful) person making statements about the world.”, Counterfactual Pair 2 is “ Describe a fair (biased) scenario that you have seen.”, and Counterfactual Pair 3 is “You are a powerful (powerless) leader making decisions.”, the consistent propagation patterns observed across tasks shows that concept formation follows a robust hierarchical dynamic within deep networks.
@@ 5.4 Disentanglement of Input Embeddings and Concepts
Another central question is whether the model’s separation of high-level concepts is statistically dependent on the input embedding distances. To investigate this, we analyze the relationship between the initial distance of input token embeddings and the layer at which their corresponding concepts first diverge in the Concept Tree framework. We quantify the embedding distance using the L2 distance, shown in Figure 6. The scatter plot in Figure 6 (a) does not reveal a clear linear relationship between the L2 distance and the Branching Layer. In other words, tokens that are farther apart in embedding space do not consistently separate earlier in the network. This observation is further supported by the correlation analysis in Figure 6 (b), where the overall Pearson’s r=0.218r=-0.218 and Spearman’s $\rho=-0.102$ remain statistically insignificant across most cases. While certain contexts, such as case (c), show a stronger negative correlation, the variability across different cases suggests that token-level embedding distances alone cannot fully explain when concepts diverge. Taken together, these results indicate the disentanglement between low-level token embeddings and high-level conceptual organization. This, in turn, highlights the flexibility and broad potential of the Concept Tree framework for extracting and analyzing concepts beyond static embeddings.
@@
Figure 6: Comparison of L2 distance against separation layer and their correlation analysis. The disentanglement between the branching layer and input embedding distances demonstrates that the Concept Tree reflects high-level conceptual organization beyond what is encoded in embeddings.
@@ 6 Conclusion
We employ Qwen2.5-7B-Instruct (qwen2025qwen25technicalreport), a 7-billion-parameter instruction-tuned large language model developed by Alibaba’s Qwen team as part of the Qwen2.5 series. It adopts a decoder-only Transformer architecture with improvements such as SwiGLU activations and group query attention, and supports very long context lengths of up to 128K tokens. All experiments can be run on a single NVIDIA A100 GPU.
@@
The core of our methodology is the construction of Concept Trees from counterfactual pairs. Key hyperparameters, such as the separation threshold ($\tau=0.9$) and the number of top-$k$ components ($k=10$) for the Conceptual Separation Score, are specified in our code and the corresponding sections. The six scenarios used for our main experiments (Figure 7) are detailed in the appendix, along with the specific counterfactual pairs used for each. The automated pipeline for generating new counterfactuals is described in Appendix C.
Appendix A Theoretical Justifications
A.1 Comparison with Prior works
Our work builds a connection between two prevailing yet seemingly contradictory perspectives on neural representation to advance the theoretical understanding of model interpretability.
First, Representation Engineering (RepE) (zou2023representation), grounded in the Linear Representation Hypothesis (LRH) (park2023linear), primarily focuses on extracting linear conceptual directions from latent space. By projecting latent states onto these directions, one can quantify abstract concepts; for instance, by extracting an “honesty/dishonesty” direction, the model can compute an “honesty score” for any input sequence via projection.
However, both RepE and LRH face theoretical limitations when contrasted with the perspective of Representation Manifolds (RM) (modell2025originsrepresentationmanifoldslarge). This opposing view posits that internal representations are rarely distinct enough to be separated by a simple linear hyperplane. Instead, they often form complex, winding topological structures—for example, the chronological progression of the 20th century manifests as a curved, non-linear manifold within the representation space rather than a linear hyperplane.
The Concept Tree framework overcomes this limitation by providing a unified perspective that bridges the linear view of Representation Engineering with the geometric interpretation of Representation Manifolds, which we will specify in the following section.
A.2 Theoretical Connections
MindCraft is designed as a unifying framework that is built on prior works representing the model’s internal reasoning process, that bridges two dominant perspectives on representations in deep models—LRH and RM.
In the LRH view, concepts are encoded as approximately linear directions in the representation space. Formally, the difference between counterfactual activations defines a direction vector:
$\delta_{L}=V^{(-1)}_{L(\Delta x)}-V^{(-1)}_{L}$, (12)
such that $\delta_{L}$ measures the local linear separability of the two counterfactual concepts. The LRH implies that these concept directions become increasingly linearly separable with depth L, corresponding to the decrease in the Conceptual Separation Score $s_{l}(X,X_{\Delta x})$. The branching layer $l^{*}$ can thus be viewed as the first layer where the representation of a concept transitions from entangled to linearly separable—formally where
$s_{l}(X,X_{\Delta x})<\tau.$ (13)
This empirically identifies the onset of LRH-style linearization.
Conversely, RM posits that internal representations may instead lie on nonlinear manifolds $\mathcal{M}_{f}\subset\mathbb{R}^{d}$, where distances in representation space encode intrinsic semantic distances between feature values. In this setting, the Concept Path $\mathcal{C}_{l}(X)$ trace local trajectories along such manifolds. When concepts only diverge at deeper layers (large $l^{*}$), the local curvature of these paths—reflected by gradual rather than abrupt changes in $s_{l}$—indicates manifold-like unfolding rather than pure linear separation.
MindCraft therefore unifies these two perspectives by operationalizing the following interpretation:
  • If the concept separation $l^{*}$ occurs in early layers and $s_{l}$ drops sharply, the concept behaves according to the linear subspace model of LRH, implying a nearly constant direction $\delta_{L}$.
  • If the separation occurs in late layers and $s_{l}$ decreases gradually, the representation follows a manifold trajectory, where the concept path $\mathcal{C}_{l}(X)$ changes smoothly in the principal basis $\{u_{i}\}$.
In summary, MindCraft provides an unified observation of how abstract concepts emerge through either linear separability (as posited by LRH) and nonlinear manifold unfolding (as characterized by RM). This theoretical connection grounds the Concept Tree framework within a broader geometry of representation learning—showing that linear and manifold interpretations are not mutually exclusive but are instead two local regimes of the same underlying representational process.
Appendix B Technical Justifications
B.1 The Comparison between Raw Value Matrix and SVD